OpenStudy (baby456):

what is a the pattern in -3,9,-27,81

OpenStudy (anonymous):

x, x^2, x^3, x^4,..........................

OpenStudy (baby456):

what does that mean

OpenStudy (raden):

looks like the geometeri sequence with ratio (r) = 9/-3 = -3 use the general formula : an = a1 * r^(n-1) given a1 = -3 and we get r = -3, therefore an = -3 * (-3)^(n-1)

OpenStudy (anonymous):

Thats the pattern it is -3, (-3)^2 , (-3)^3, (-3)^4, ............

OpenStudy (anonymous):

pattern is -3